OpenStudy (anonymous):

a solid is formed by rotating y =e^x +1 between x=0 and x=3. Find the volume of the solid formed.

OpenStudy (anonymous):

rotating about what

OpenStudy (anonymous):

x axis ?

OpenStudy (anonymous):

most likely, but can you have to specify

OpenStudy (anonymous):

i suppose but question does not mention anything

OpenStudy (anonymous):

\[= \pi \int\limits_{0}^{3} ( e^x +1)^2 dx = \pi \int\limits_{0}^{3} ( e^{2x} +2e^x +1 ) dx\]

OpenStudy (anonymous):

\[= \pi [ \frac{1}{2} e^{2x}+ 2e^x +x ]\] between the limits

OpenStudy (anonymous):

so what did you get as your answer?

OpenStudy (anonymous):

\[= \pi [ ( \frac{1}{2} e^6 + 2e^3 + 3 ) - ( \frac{1}{2} +2 +0) ] = \pi [ \frac{1}{2} e^6 + 2e^3 +\frac{1}{2}] \]

OpenStudy (anonymous):

\[= \frac{\pi}{2} [ e^6 +4e^3 +1] \]

OpenStudy (anonymous):

i got the same answer as you but the answer provided says otherwise. it says (\[\pi/2 (e^2+1)^2u^2\]

OpenStudy (anonymous):

any ideas on how they got that answer?

OpenStudy (anonymous):

no, fairly sure they are wrong.

OpenStudy (anonymous):

i'm curious about the u. where does that come from? and i don't think it is a typo because the next questions asks "using the substitution u=logex, find \[\int\limits_{1}^{e} loge/x .dx\]